OpenStudy (anonymous):

Please help!!!!!!!!!!!!! Use basic identities to simplify the expression. sin2θ + tan2θ + cos2θ

OpenStudy (anonymous):

would the answer be \[\sec^2\theta \]

OpenStudy (anonymous):

@Mertsj ?

OpenStudy (mertsj):

\[\sin 2\theta +\frac{\sin 2\theta}{\cos 2\theta}+ \cos 2\theta=\] \[2\sin \theta \cos \theta+\frac{2\sin \theta \cos \theta}{\cos ^2\theta-\sin ^2\theta}+\cos ^2\theta-\sin ^2\]

OpenStudy (mertsj):

I guess I'm not seeing how you get sec^2 theta.

OpenStudy (mertsj):

What happened to all the double angles?

OpenStudy (anonymous):

i didnt mean sec i meant sin my bad

OpenStudy (mertsj):

So you got this for the answer: \[\sin ^2\theta \]

OpenStudy (anonymous):

just sin theta

OpenStudy (mertsj):

Could you post your work? I can't see how you got rid of sin(2theta) and tan(2theta) and cos(2theta) without using the double angle identities.

OpenStudy (mertsj):

Oh gees. It just hit me...Did you mean sin^2 theta and cos^2 theta and tan^2 theta instead of sin(2theta)...

OpenStudy (anonymous):

yes i did sorry i didnt clarify

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

When you want to indicate an exponent use ^. sec^2 theta is the correct answer.

OpenStudy (anonymous):

oh okay thank you so much! I will always clarify from now on lol